# Explore Synergistic Interaction Across Frames for Interactive Video Object Segmentation

### v1

@@ III-C Across-Round Propagation Module
where $\mathcal{A}(\cdot)$ denotes the Across-Round Memory Update network (MU-Net, adopted AOT). ${U}^{r}$ denotes the across-round memory in the $r$-th interaction round, $I_{t}\in\mathbb{R}^{THW\times C}$ is the RGB image of the interacted frame, and $M_{{t}}^{r}\in\mathbb{R}^{THW\times C}$ is the mask of the interacted frames with identification tags.
Propagation with Batch Decoder. We put the image and the across-frame memory into the ID-queried batch decoder on a frame-by-frame basis to generate the corresponding multi-object mask. The batch decoder contains multiple object-level channels that can simultaneously process multiple objects with identification tags,
@@
where $\mathcal{P}(\cdot)$ denotes the batch decoder. $I_{t}\in\mathbb{R}^{THW\times C}$ denotes the RGB image of $t$-th frame that without interaction. ${\bf U}^{r}$ denotes the across-frame memory and $M_{t}^{r}\in\mathbb{R}^{THW\times C}$ denotes the output mask of $t$-th non-interactive frame. $\widetilde{\bf R}$ is the set of frames without interaction.
@@ IV Implementation Details
Datasets. Our model is trained on the Lvis [43] (static image dataset), YouTubeVOS (YTB) [44] and DAVIS 2017 [45]. The three datasets vary in quantity and quality of annotations: 1) Lvis is designed for large vocabulary instance segmentation, containing instance masks for over 1000 entry-level object categories. We obtained multiple images from the static images in Lvis by applying translations, rotations, and other morphological changes. We then treated these images as a video and conducted pre-training for SIAF using this modified dataset. 2) YouTube-VOS is the latest large-scale benchmark for multi-object video segmentation. Specifically, YouTube-VOS contains 3471 videos in the training split with 65 categories and 474/507 videos in the validation 2018/2019 split with an additional 26 unseen categories. The unseen categories do not exist in the training split to evaluate the algorithms’ generalization ability. 3) DAVIS 2017 contains 60 training sequences and 30 validate sequences. We use the DAVIS 2017 training set, which provides high-quality densely annotated segmentation mask annotation for each frame. For evaluation, we employ 30 validate sequences.
Figure 6: The inference speed varies with the number of objects (§V-A). The time consumption of MiVOS is increased with the number of objects, but SIAF processes multiple objects as efficiently as a single object. Note that we set the batch in the decoder as 10, so the time consumed increases as the object number is greater than 10, and we have the flexibility to train models with different batch numbers as needed.
To evaluate the effectiveness of SIAF, we conducted experiments on the DAVIS 2017. Currently, DAVIS is the sole benchmark for systematically validating iVOS. It offers diverse scribble types, automatic scribble generation, and an interactive evaluation environment. Thus, we follow previous SOTA work to develop SIAF on it.
Scribbles Generation. The robots in DAVIS can provide scribbles for a selected frame. We generated scribbles for YouTube-VOS consistent with the DAVIS methodology. To ensure fairness during inference, we adopt the DAVIS test setting. Specifically, we use three initial scribbled frames provided by DAVIS, and use the toolkit to select the worst predicted frame and scribble it. Further details about datasets can be found in the supplementary materials.
@@
Training Details. We implement SIAF by PyTorch [46], and train all models on a 24GB GPU device (RTX 3090Ti). When training on the YouTube-VOS and DAVIS datasets, we employ the AdamW optimizer [47] with a learning rate of $2\times 10^{-4}$ and $2\times 10^{-5}$, respectively. Our batch size is set to 4. The AFI Module is first trained on synthetic video sequences by Lvis, then transferred to YouTube-VOS and DAVIS 2017. The loss function is a 0.5:0.5 combination of bootstrapped cross-entropy loss and soft Jaccard loss [48].
@@
R50-SIAF 85.2 87.4
SwinB-SIAF 86.2 88.2
R50-SIAF (+Y) 85.4 87.3
${}^{\dagger}$ (+Y) 86.2 88.7
@@
R50-SIAF 86.9 88.6
SwinB-SIAF 87.5 89.3
R50-SIAF (+Y) 87.0 89.1
SwinB-SIAF (+Y) 87.5 89.6
V Experiments
V-A Comparison at Speed.
Settings. In Fig. 6, we compare the inference speed of SIAF with the state-of-the-art model MiVOS, across varying numbers of target objects. The maximum number of objects SIAF can concurrently handle is dictated by the channel quantity established during training, set in this instance to accommodate up to ten objects simultaneously. Consequently, we present an experiment of inference speed, from processing a single object to concurrently handling eleven objects, clearly illustrating the model’s inference speed modulation in response to increasing object quantity.
TABLE II: The ablation study on the DAVIS with multi-frame and multi-round scenario ($f=3$, $r=3$) (§V-C). We first trained SIAF on Lvis and we took a network that without Across-Frame Attention and Re-propagation strategy as the baseline.
SwinB-SIAF $\mathcal{J\&F}@60$
@@ V Experiments
Results. SIAF has an excellent performance in terms of speed. From the results in Fig. 6, we get the following findings: 1) MiVOS exhibits a progressive rise in inference time correlating with the increase in target objects (from one object at 9.25 FLOP to ten objects at 95.2 FLOPS). In contrast, our SIAF consistently maintains a steady inference speed of 11.95 FLOPS even in scenarios involving ten objects. Notably, this implies that SIAF can run more than 3× faster than the current state-of-the-art model under challenging multi-object multi-frame scenarios. 2) A notable uptick in inference time for our SIAF model becomes evident when processing the eleventh object. This phenomenon can be attributed to the design of our trained decoder, which operates optimally with a specific batch number of 10, enabling simultaneous processing of up to 10 objects. Consequently, when handling 11 objects, there is an inevitable increase in processing time. To cater to diverse needs, we maintain flexibility in training models with varying batch numbers as required.
@@
Settings. SIAF can process multi-frame scribble input, but previous methods could only process single-frame scribble input. To ensure fairness, we compared SIAF with competitors under the single-frame setting ($f=1$). Additionally, to validate the effectiveness of our multi-object multi-frame interaction, we also provide the performance of SIAF under the new setting of multiple frames ($f=3$), see Tab. I.
@@
SIAF introduces a novel capability: processing multi-frame interactive information within a single iteration, a feature absent in previous methods for single iterations. To evaluate this, we conducted experiments, comparing our model against established approaches and the current state-of-the-art method. The assessment involved a dual-setting analysis: 1) Old Setting: Robots could annotate only one frame per iteration and propagation. Following annotation, they received the segmentation result after the initial propagation round. Subsequently, they adjusted annotations for a specific frame, proceeding to the second iteration and propagation, culminating in the segmentation result after the second round. Tab. I presents segmentation results for the first and third rounds in the single-frame scenario. To maintain parity, our SIAF model underwent training in the single-frame datasets. 2) New Setting: Robots could annotate multiple frames per iteration and propagation. After annotating multiple frames, they received the segmentation result after the initial propagation. Subsequently, robots selected multiple frames from the segmentation result for interaction and modification, yielding the segmentation result after the second propagation round. We showcased the segmentation results for the first and third rounds under this new setting, where interactions involved three frames per iteration. The term ’robots’ here signifies annotations sourced from the DAVIS 2017 dataset, capable of providing annotations for a selected frame.
@@
Results. Tab. I shows that SIAF achieves new state-of-the-art performance (89.6% $\mathcal{J\&F}@60$) when across-frame scribbles are available. To ensure fairness, our SIAF was trained under a single-frame scenario when we compared it to the previous single-frame input methods. In a word, SIAF can achieve the SOTA performance regardless of the setting of single-frame/multi-frame and single-round/multi-round.
@@
From the results in Tab. I, we find that: 1) Despite notable improvements in inference speed, SIAF consistently outperforms the state-of-the-art model. In the ’old setting’, compared to the MiVOS (85.5 J&F@60 in the first round and 88.5 J&F@60 in the third round), our SwinB-SIAF demonstrates superior performance (86.2 J&F@60 in the first round and 88.7 J&F@60 in the third round). 2) In the ’new setting’, SwinB-SIAF achieves peak performance (87.5 J&F@60 in the first round and 89.6 J&F@60 in the third round). Notably, SIAF is the pioneering model capable of assimilating multi-frame interactive information within a single iteration. This feature not only enriches user interaction but also bolsters processing speed and model efficacy after a single iteration. 3) Results from multiple iterations surpass those from a single iteration. This can be attributed to users iteratively refining segmentation via scribbles or click-based interactions before each iteration. More interactions lead to segmentation aligning more closely with user-defined objectives.
@@
R50 SIAF $\mathcal{J\&F}@60$ SwinB SIAF $\mathcal{J\&F}@60$
@@
R50-SIAF Memory(MB) Time
@@
Figure 7: The interactive graphical user interface (GUI; §V-E). In (a), a succinct GUI streamlines user operations, enhancing the user experience. In (b), SIAF entails the creation of masks for multi-object and multi-frame datasets. In (c), our model exhibits commendable efficacy even in the case of the wild.
@@ V-C Ablation Study
Results. From the results in Tab. II and Tab. III, we find that: 1) Effectiveness of Across-Frame Interaction. The Across-Frame Interaction Module (AFI) can accept multiple frames and generate their masks. Moreover, the AFI incorporates an Across-Frame Attention (AFA) that effectively aggregates multi-scale semantic information and captures temporal information across the interacted frames (see Fig. 4). Tab. II demonstrates that multi-scale semantic and temporal information can improve performance. 2) Effectiveness of Across-Round Propagation. Multi-round interactions have been demonstrated to outperform single-round interactions because each interaction round contains more user intent. Thus, we design the across-round memory to store important interaction information and re-propagate it for all frames without interaction. Tab. II shows the benefit of this propagation strategy. 3) Performance Varies with Frame Number and Interaction Rounds. In Tab. III, as the number of interacting frames increases within the same iteration round, users can furnish more guiding information, consequently enhancing SIAF performance. For instance, SwinB-SIAF demonstrates 86.2 at $J\&F$ in a single round with individual frames, escalating to $87.5$ at $J\&F$ in a round involving three frames. Hence, our proposed SIAF, facilitating multiple user interactions within a single round, proves indispensable. This approach not only grants users greater interaction flexibility but also enhances accuracy within identical iteration rounds. Moreover, as the iteration count rises, meticulous tuning of each iteration’s rounds based on previous segmentation augments the model’s performance.
Figure 8: Qualitative comparison on the DAVIS interactive track (§V-B). We compare our method SIAF with MiVOS [6] on the DAVIS dataset.
@@ V-D Memory and Time Consumed Analysis
3) Memory and Time Consumed for Each Subcomponent. Tab. IV illustrates the memory and time demands of submodules. Notably, within the across-round propagation module, MU-Net and MP-Net dominate memory usage. In contrast, our proposed multi-frame interaction module maintains a manageable memory. Moreover, the propagation module necessitates more inference time compared to the interaction module. Overall, SIAF exhibits commendable and controlled performance in both memory usage and inference time.
@@ V-E Interactive Graphical User Interface
We developed an interactive graphical user interface (GUI) that enables multi-frame interaction for mask generation. In Fig. 7 (a), a succinct visual interactive system streamlines user operations, enhancing the user experience. The graphical user interface features intuitive buttons for sketching annotations and generating corresponding frame segmentation results. Additionally, a dedicated button facilitates the propagation of segmentation results from interacted frames to the entirety of the sequence. Moving to Fig. 7 (b), the procedural framework entails the creation of masks for multi-object and multi-frame datasets. Initially, users designate a frame for annotating scribbles, thereby enabling the annotation of multiple objects within the frame. Subsequently, users have the option to extend their annotations to additional frames. Upon completion, activating the ”commit scribbles” button produces segmentation results encompassing all annotated frames. A concluding step involves users clicking the ”propagate” button, yielding comprehensive segmentation results for the entire video. In Fig. 7 (c), our model exhibits commendable efficacy even in the case of the wild.
Using the interactive GUI tool, users can annotate multiple objects on multiple frames, and the SIAF automatically generates masks for all frames concurrently. If the user finds the generated mask to be unsatisfactory, another round of interaction can be carried out to improve it. SIAF then propagates the refined masks to all frames in the video, resulting in the segmentation of the entire video. SIAF enables efficient and accurate mask generation while reducing the user’s workload and time spent on propagation.

### v3

@@ III-C Concurrent Propagation Module
where $\mathcal{A}(\cdot)$ denotes the Concurrent Memory Update network (MU-Net, adopted AOT). ${U}^{r}$ denotes the across-round memory in the $r$-th interaction round, $I_{t}\in\mathbb{R}^{THW\times C}$ is the RGB image of the interacted frame, and $M_{{t}}^{r}\in\mathbb{R}^{THW\times C}$ is the mask of the interacted frames with identification tags.
Propagation with ID-queried Decoder. We put the image and the across-frame memory into the ID-queried decoder on a frame-by-frame basis to generate the corresponding multi-object mask. The ID-queried decoder contains multiple object-level channels that can simultaneously process multiple objects with identification tags,
@@
where $\mathcal{P}(\cdot)$ denotes the ID-queried decoder. $I_{t}\in\mathbb{R}^{THW\times C}$ denotes the RGB image of $t$-th frame that without interaction. ${\bf U}^{r}$ denotes the across-frame memory and $M_{t}^{r}\in\mathbb{R}^{THW\times C}$ denotes the output mask of $t$-th non-interactive frame. $\widetilde{\bf R}$ is the set of frames without interaction.
@@ IV Implementation Details
Datasets. Our model is trained on the Lvis [87] (static image dataset), YouTubeVOS (YTB) [88] and DAVIS 2017 [89]. The three datasets vary in quantity and quality of annotations: 1) Lvis is designed for large vocabulary instance segmentation, containing instance masks for over 1000 entry-level object categories. We obtained multiple images from the static images in Lvis by applying translations, rotations, and other morphological changes. We then treated these images as a video and conducted pre-training for IDPro using this modified dataset. 2) YouTube-VOS is the latest large-scale benchmark for multi-object video segmentation. Specifically, YouTube-VOS contains 3471 videos in the training split with 65 categories and 474/507 videos in the validation 2018/2019 split with an additional 26 unseen categories. The unseen categories do not exist in the training split to evaluate the algorithms’ generalization ability. 3) DAVIS 2017 contains 60 training sequences and 30 validate sequences. We use the DAVIS 2017 training set, which provides high-quality densely annotated segmentation mask annotation for each frame. For evaluation, we employ 30 validate sequences.
Figure 6: The computational complexity varies with the number of objects (§V-A). “FLOPS” stands for “Floating Point Operations Per Second” and is a measure of the computational complexity of a system. The computational complexity of MiVOS is increased with the number of objects, but IDPro processes multiple objects as efficiently as a single object. Note that we set the batch in the decoder as 10, so the time consumed increases as the object number is greater than 10, and we have the flexibility to train models with different batch numbers as needed.
Figure 7: The inference speed varies with the number of objects (§V-A). The time consumption of MiVOS is increased with the number of objects, but IDPro processes multiple objects as efficiently as a single object.
To evaluate the effectiveness of IDPro, we conducted experiments on the DAVIS 2017. Currently, DAVIS is the sole benchmark for systematically validating iVOS. It offers diverse scribble types, automatic scribble generation, and an interactive evaluation environment. Thus, we follow previous SOTA work to develop IDPro on it.
Scribbles Generation. The robots in DAVIS can provide scribbles for a selected frame. We generated scribbles for YouTube-VOS (YTB) consistent with the DAVIS methodology. To ensure fairness during inference, we adopt the DAVIS test setting. Specifically, we use three initial scribbled frames provided by DAVIS, and use the toolkit to select the worst predicted frame and scribble it.
@@
Training Details. We implement IDPro by PyTorch, and train all models on a 24GB GPU device (RTX 3090Ti). When training on the YouTube-VOS and DAVIS datasets, we employ the AdamW optimizer [90] with a learning rate of $2\times 10^{-4}$ and $2\times 10^{-5}$, respectively. Our batch size is set to 4. The AFI Module is first trained on synthetic video sequences by Lvis, then transferred to YouTube-VOS and DAVIS 2017. The loss function is a 0.5:0.5 combination of bootstrapped cross-entropy loss and soft Jaccard loss [91].
V Experiments
V-A Comparison at Speed.
Settings. In Fig. 7, we compare the inference speed of IDPro with the state-of-the-art model MiVOS, across varying numbers of target objects. The maximum number of objects IDPro can concurrently handle is dictated by the channel quantity established during training, set in this instance to accommodate up to ten objects simultaneously. Consequently, we present an experiment of inference speed, from processing a single object to concurrently handling eleven objects, clearly illustrating the model’s inference speed modulation in response to increasing object quantity.
@@ V Experiments
R50-IDPro 85.2 87.4
SwinB-IDPro 86.2 88.2
R50-IDPro (+Y) 85.4 87.3
SwinB-IDPro (+Y) 86.2 88.7
@@
R50-IDPro 86.9 88.6
SwinB-IDPro 87.5 89.3
R50-IDPro (+Y) 87.0 89.1
SwinB-IDPro (+Y) 87.5 89.6
TABLE II: The ablation study on the DAVIS with multi-frame and multi-round scenario ($f=3$, $r=3$) (§V-C). We first trained IDPro on Lvis and we took a network that without Across-Frame Attention and Re-propagation strategy as the baseline.
SwinB-IDPro $\mathcal{J\&F}@60$
@@
Results. IDPro has an excellent performance in terms of speed. From the results in Fig. 7, we get the following findings: 1) MiVOS exhibits a progressive rise in inference time correlating with the increase in target objects (from one object at 9.25 FLOP to ten objects at 95.2 FLOPS). In contrast, our IDPro consistently maintains a steady inference speed of 11.95 FLOPS even in scenarios involving ten objects. Notably, this implies that IDPro can run more than 3× faster than the current state-of-the-art model under challenging multi-object multi-frame scenarios. 2) A notable uptick in inference time for our IDPro model becomes evident when processing the eleventh object. This phenomenon can be attributed to the design of our trained decoder, which operates optimally with a specific batch number of 10, enabling simultaneous processing of up to 10 objects. Consequently, when handling 11 objects, there is an inevitable increase in processing time. To cater to diverse needs, we maintain flexibility in training models with varying batch numbers as required.
@@
Settings. IDPro can process multi-frame scribble input, but previous methods could only process single-frame scribble input. To ensure fairness, we compared IDPro with competitors under the single-frame setting ($f=1$). Additionally, to validate the effectiveness of our multi-object multi-frame interaction, we also provide the performance of IDPro under the new setting of multiple frames ($f=3$), see Tab. I.
@@
IDPro introduces a novel capability: processing multi-frame interactive information within a single iteration, a feature absent in previous methods for single iterations. To evaluate this, we conducted experiments, comparing our model against established approaches and the current state-of-the-art method. The assessment involved a dual-setting analysis: 1) Old Setting: Robots could annotate only one frame per iteration and propagation. Following annotation, they received the segmentation result after the initial propagation round. Subsequently, they adjusted annotations for a specific frame, proceeding to the second iteration and propagation, culminating in the segmentation result after the second round. Tab. I presents segmentation results for the first and third rounds in the single-frame scenario. To maintain parity, our IDPro model underwent training in the single-frame datasets. 2) New Setting: Robots could annotate multiple frames per iteration and propagation. After annotating multiple frames, they received the segmentation result after the initial propagation. Subsequently, robots selected multiple frames from the segmentation result for interaction and modification, yielding the segmentation result after the second propagation round. We showcased the segmentation results for the first and third rounds under this new setting, where interactions involved three frames per iteration. The term ’robots’ here signifies annotations sourced from the DAVIS 2017 dataset, capable of providing annotations for a selected frame.
@@
Results. Tab. I shows that IDPro achieves new state-of-the-art performance (89.6% $\mathcal{J\&F}@60$) when across-frame scribbles are available. To ensure fairness, our IDPro was trained under a single-frame scenario when we compared it to the previous single-frame input methods. In a word, IDPro can achieve the SOTA performance regardless of the setting of single-frame/multi-frame and single-round/multi-round.
@@
From the results in Tab. I, we find that: 1) Despite notable improvements in inference speed, IDPro consistently outperforms the state-of-the-art model. In the ’old setting’, compared to the MiVOS (85.5 $J\&F@60$ in the first round and 88.5 $J\&F@60$ in the third round), our SwinB-IDPro demonstrates superior performance (86.2 $J\&F@60$ in the first round and 88.7 $J\&F@60$ in the third round). 2) In the ’new setting’, SwinB-IDPro achieves peak performance (87.5 $J\&F@60$ in the first round and 89.6 $J\&F@60$ in the third round). Notably, IDPro is the pioneering model capable of assimilating multi-frame interactive information within a single iteration. This feature not only enriches user interaction but also bolsters processing speed and model efficacy after a single iteration. 3) Results from multiple iterations surpass those from a single iteration. This can be attributed to users iteratively refining segmentation via scribbles or click-based interactions before each iteration. More interactions lead to segmentation aligning more closely with user-defined objectives.
@@
R50-IDPro $\mathcal{J\&F}@60$ SwinB-IDPro $\mathcal{J\&F}@60$
@@
R50-IDPro Memory(MB) Time
@@
TABLE VII: Sensitivity Study of Scribbles (§V-C). We divided scribbles into three qualities: detailed, general, and rough to verify their performance separately.
Scribble MiVOS IDPro
Detailed 85.9 86.3
General 85.1 85.2
Rough 84.8 85.0
Figure 8: The interactive graphical user interface (GUI; §V-E). In (a), a succinct GUI streamlines user operations, enhancing the user experience. In (b), IDPro entails the creation of masks for multi-object and multi-frame datasets. In (c), our model exhibits commendable efficacy even in the case of the wild.
@@ V-C Ablation Study
Results. From the results in Tab. II and Tab. III, we find that:
1) Effectiveness of Across-Frame Interaction. The Across-Frame Interaction Module (AFI) can accept multiple frames and generate their masks. Moreover, the AFI incorporates an Across-Frame Attention (AFA) that effectively aggregates multi-scale semantic information and captures temporal information across the interacted frames (see Fig. 4). Tab. II demonstrates that multi-scale semantic and temporal information can improve performance.
2) Effectiveness of Concurrent Propagation. Multi-round interactions have been demonstrated to outperform single-round interactions because each interaction round contains more user intent. Thus, we design the across-round memory to store important interaction information and re-propagate it for all frames without interaction. Tab. II shows the benefit of this propagation strategy.
3) Performance Varies with Frame Number and Interaction Rounds. In Tab. III, as the number of interacting frames increases within the same iteration round, users can furnish more guiding information, consequently enhancing IDPro performance. For instance, SwinB-IDPro demonstrates 86.2 at $J\&F$ in a single round with single frames, escalating to $87.5$ at $J\&F$ in a round involving three frames. Hence, our proposed IDPro, facilitating multiple user interactions within a single round, proves indispensable. This approach not only grants users greater interaction flexibility but also enhances accuracy within identical iteration rounds. Moreover, as the iteration count rises, meticulous tuning of each iteration’s rounds based on previous segmentation augments the model’s performance.
4) Performance Varies with the level of detail scribbles.We divided scribbles into three qualities: detailed, general, and rough following. Specifically, IDPro achieves 86.3, 85.2 (-1.1), 85.0 (-1.3) ($\mathcal{J\&F}@60$) on the three scribble types respectively. Compared to SOTA, which achieves 85.9, 85.1 (-0.8), and 84.8 (-1.1). Experimental results yield that the level of detail in user-provided scribbles significantly impacts performance. Specifically, more intricacy within the scribbles corresponds to improved output quality. Moreover, the consistent efficacy of IDPro across varying degrees of meticulous scribbling. These observations underscore the pivotal role of scribble precision in influencing outcomes and emphasize the stability of IDPro performance.
Figure 9: Qualitative comparison on the DAVIS interactive track (§V-B). We compare our method IDPro with MiVOS [6] on the DAVIS dataset.
@@ V-D Memory and Time Consumed Analysis
3) Memory and Time Consumed for Each Subcomponent. Tab. IV illustrates the memory and time demands of submodules. Notably, within the concurrent propagation module, MU-Net and MP-Net dominate memory usage. In contrast, our proposed multi-frame interaction module maintains a manageable memory. Moreover, the propagation module necessitates more inference time compared to the interaction module. Overall, IDPro exhibits commendable and controlled performance in both memory usage and inference time.
@@ V-E Interactive Graphical User Interface
We developed an interactive graphical user interface (GUI) that enables multi-frame interaction for mask generation. In Fig. 8 (a), a succinct visual interactive system streamlines user operations, enhancing the user experience. The graphical user interface features intuitive buttons for sketching annotations and generating corresponding frame segmentation results. Additionally, a dedicated button facilitates the propagation of segmentation results from interacted frames to the entirety of the sequence. Moving to Fig. 8 (b), the procedural framework entails the creation of masks for multi-object and multi-frame datasets. Initially, users designate a frame for annotating scribbles, thereby enabling the annotation of multiple objects within the frame. Subsequently, users have the option to extend their annotations to additional frames. Upon completion, activating the “commit scribbles” button produces segmentation results encompassing all annotated frames. A concluding step involves users clicking the “propagate” button, yielding comprehensive segmentation results for the entire video. In Fig. 8 (c), our model exhibits commendable efficacy even in the case of the wild.
Using the interactive GUI tool, users can annotate multiple objects on multiple frames, and the IDPro automatically generates masks for all frames concurrently. If the user finds the generated mask to be unsatisfactory, another round of interaction can be carried out to improve it. IDPro then propagates the refined masks to all frames in the video, resulting in the segmentation of the entire video. IDPro enables efficient and accurate mask generation while reducing the user’s workload and time spent on propagation.